### v2

@@ 2.4   Interpreting the parallel trends assumption in the outcome
So far, we have formalized the target parameter and identifying assumptions in $2\times 2$ DID-IV designs. Overall, in this design, we rely on the common DID identification strategy in both the denominator and the numerator of the Wald-DID estimand in order to identify the effects of the instrument on the treatment and the outcome. The interpretation of the parallel trends assumption in the outcome, however, is less clear than that of the parallel trends assumption in the canonical DID designs because it does not depend on the untreated outcome, but depends on the unexposed outcome.
@@
In this section, we clarify the interpretation of the parallel trends assumption in the outcome in $2\times 2$ DID-IV designs. To do so, we first uncover the selection mechanism behind the canonical DID-IV designs. Specifically, we show that in $2\times 2$ DID-IV designs, time also plays the role of instrument, and investigate how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. We then show that the average time trend in the outcome is equal to a weighted average of the indirect effect of time on the outcome through treatment. We call this an expected time gain and show that the parallel trends assumption in the outcome requires the homogeneous expected time gain between exposed and unexposed groups.
@@
First, we introduce the additional notation. Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$, which represents the treatment adoption process from period 00 to period 1 if the instrument path were $z=(0,0)$: the first element $D_{i,0}((0,0))$ represents the potential treatment choice without instrument in period 00 and the second element $D_{i,1}((0,0))$ represents the potential treatment choice without instrument in period $1$.
@@
We recall that in $2\times 2$ DID-IV designs, the units are allowed to take the treatment without instrument during the two periods. In this setting, one can see that time also plays the role of instrument, that is, time also affects one’s treatment choices during the two periods.
@@
Similar to the group variable $G_{i}^{Z}$ introduced in section 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ to be the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ to be the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ to be the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ to be the time defiers. The time never-takers $NT^{T}$ and the time always-takers $AT^{T}$ do not change their treatment status without instrument during the two periods. In other words, their treatment choices are time-invariant in the absence of the policy shock (instrument). On the other hand, the time compliers $CM^{T}$ are induced to treatment in period $1$ without instrument, and the time defiers $DF^{T}$ leave the treatment in period $1$ without instrument, that is, their treatment choices are affected by time in the absence of the policy shock.
@@
By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we can describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs. Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, the types painted in gray color in both tables.
@@
In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but can not observe $D_{1}((0,0))$ for each unit (see Table 1). Here, we use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$. In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but can not observe $D_{1}((0,1))$ for each unit (see Table 2). If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time. If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument.
@@
Given these tables, we now consider the interpretation of the parallel trends assumption in the outcome. Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in exposed and unexposed groups, respectively:
@@
We note that the parallel trends assumption in the outcome requires $\Delta_{0}=\Delta_{1}$. At the same time, we also note that $\Delta_{0}$ is estimable by the observed change in outcomes in an unexposed group during the two periods, whereas $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
To make the interpretations of $\Delta_{0}$ and $\Delta_{1}$ clear, we introduce the additional notation. For each time group $g^{t}\in\{AT^{T},NT^{T},CM^{T},DF^{T}\}$, let $w_{g^{t},e}\equiv Pr[G_{i}^{T}=g^{t}|E_{i}=e]$ denote the population share in group $E_{i}=e\in\{0,1\}$. Let $\Delta_{g^{t},e}$ denote the expectation of the potential outcome path in the absence of the policy shock for time group $g^{t}$ in group $E_{i}=e$:
@@
For instance, $w_{CM^{T},0}=Pr[CM^{T}|E_{i}=0]$ and $\Delta_{CM^{T},0}=E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=0]$ represent the population share and the expectation of the outcome path without instrument for the time compliers in an unexposed group, respectively.
Each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are estimable in an unexposed group ($e=0$): we can identify the time group of unit $i$ in an unexposed group and can observe the evolution of the outcome without instrument. For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$ for unit $i$ in an unexposed group, we can figure out that unit $i$ is the time compliers $CM^{T}$, and can observe the potential outcome path $(Y_{i,0}(0),Y_{i,1}(1))$ without instrument for that unit. On the other hand, each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are unobservables in an exposed group ($e=1$) because we can not specify the time group of unit $i$ in that group. For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$ for unit $i$ in an exposed group, we can not distinguish whether unit $i$ belongs to the time compliers $CM^{T}$ or the time never takers $NT^{T}$ (see Table 1).
We can interpret each $\Delta_{g^{t},e}$ as the indirect effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$. First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes respectively:
@@
This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods. If the exclusion restriction of time on the outcome holds, that is, $Y_{i,1}(1)=Y_{i,0}(1)$ and $Y_{i,1}(0)=Y_{i,0}(0)$ hold for all units $i$ (in other words, the stationary conditions hold), these quantities are equal to zero. Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms:
@@
where the weight $w_{g^{t},e}$ is non-negative and sum to one, that is, $\sum_{g^{t}}w_{g^{t},e}=1$ holds.
@@
In light of the view discussed so far, we can conclude that the parallel trends assumption in the outcome (Assumption 6), i.e., $\Delta_{0}=\Delta_{1}$ requires the same expected time gain between exposed and unexposed groups, that is, the whole indirect effects of time on the outcome through treatment are the same on average between the two groups.
Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the indirect effect of the instrument on the outcome through treatment from that of time in an exposed group. The underlying idea to capture the former effect is parallel to that of DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period. If there exist pre-periods before the exposure to the instrument, we can indirectly check the validity of Assumption 6 as well as Assumption 5, comparing the evolution of the outcome between exposed and unexposed groups. The formal discussion for the pretrend test is provided in section 5.3.
@@ 2.5   Relationship between DID and DID-IV
In this section, we describe the relationship between DID and DID-IV designs, and provide the view that DID-IV is DID with the noncompliance of the assigned treatment.
To see this, we first introduce the notation in DID designs according to section 2.1. Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program, which distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$. In the canonical DID set up, $Z_{i}=(Z_{i,0},Z_{i,1})$ takes only two values: $Z_{i}=(0,1)$ if unit $i$ is assigned to the program in period $1$ (treatment group) and $Z_{i}=(0,0)$ if unit $i$ is not assigned to the program in period $1$ (control group). Let $E_{i}\in\{0,1\}$ denote the group variable: $E_{i}=1$ if unit $i$ belongs to the treatment group and $E_{i}=0$ if unit $i$ belongs to the control group. Hereafter, in this section, we refer to treatment and control groups as exposed and unexposed groups, respectively.
We note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units. First, we assume that all the units are not allowed to receive the treatment in period 00:
@@
Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and should not receive the treatment in period $1$ if that unit belongs to an unexposed group:
@@
These restrictions imply that we have Tables 3-4 in DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
@@
Second, the ATT, the target parameter in DID designs, is equal to the LATET in the canonical DID set up. This follows from Tables 3 - 4:
@@
We note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$). We therefore define the LATET as our target parameter in DID-IV designs in section 2.2.
Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups. From Table 3 and Table 4, we can rewrite the parallel trends assumption in the untreated outcome as follows:
@@
where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively in DID designs:
@@
The above discussions clarify that DID-IV is a natural generalization of DID, and the difference between the two designs arises because DID-IV allows the noncompliance of the assigned treatment: $D_{i,t}\neq Z_{i,t}$. In the next section, we compare DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018).
@@ 2.6   Comparing DID-IV to Fuzzy DID
de Chaisemartin and D’Haultfœuille (2018) (Henceforth, ”dCDH” in this section) also investigate the identification assumptions for the Wald-DID estimand to capture the causal effects under the same setting in this article. However, they formalize $2\times 2$ DID-IV designs differently, and call them Fuzzy DID. In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID. The detailed discussion is contained in Appendix E, and this section provides the main points of that discussion.
The main difference between dCDH and this article is in the definition of target parameter: we focus on the LATET, whereas dCDH focus on the SLATET (see section 2.2 in this article). Accordingly, the identifying assumptions in Fuzzy DID designs are different from those in $2\times 2$ DID-IV designs. Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
In Appendix E, we first show that dCDH’s identifying assumptions ex ante impose strong and asymmetric restrictions on the treatment adoption behavior across units between exposed and unexposed groups. Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Table 5 and Table 6333Because dCDH consider repeated cross section settings, we can observe either $D_{0}(0)$ or $D_{1}(1)$ for each unit in Tables 5-6.. These tables are in line with Table 1 and Table 2 in this article, and clarify the difference between dCDH’s restrictions and ours: we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by monotonicity assumption, whereas dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group. This difference is ciritical for empirical application: in contrast to the monotonicity assumption, dCDH’s additional restrictions can be difficult to defend from the institutional knowledge of the problem in a given application.
@@
Next, under Tables 5-6, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters. Specifically, we show that the SLATET is a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $CM^{Z}\land AT^{T}$, which is the sub-population of the compliers $CM^{Z}$, in an exposed group. This decomposition result has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET. Second, this parameter may not be policy-relevant even when the researchers use the policy change of their interest as an instrument for the treatment because it generally contains the treatment effects among the time compliers $CM^{T}$ who are affected by time but not affected by the instrument.
@@
Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided. In this discussion, we also clarify why dCDH ex ante exclude so many types in exposed and unexposed groups respectively for the Wald-DID estimand to identify the SLATET.
@@
First, we show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome in DID designs. Recall that in DID-IV designs, we do not assume the parallel trends assumption in the untreated outcome, but assume the parallel trends assumption in the unexposed outcome. Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption generally fails to capture the average time trend of the outcome in an unexposed group. We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET. In the proof of this argument, we also demonstrate that the strong restrictions imposed by dCDH on treatment adoption behavior is also induced by imposing the parallel trends assumption in the untreated outcome.
@@
Next, we argue that the parallel trends assumption in the untreated outcome is unsuitable for DID-IV settings for three reasons. First, as already noted, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group. Second, in most applications of the DID-IV method, we can not impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al. (2005)). For instance, in Black et al. (2005), this assumption requires that the data contain the parents with zero education attainment, which is unrealistic in practice. Finally, in DID-IV settings, we can not generally test this assumption using pre-exposed period data, as some units can adopt the treatment before period 00. Indeed, we show that the placebo test proposed by dCDH is incomplete for assessing the validity of this assumption. Surprisingly, we also show that their placebo test coincides with our pretrends test, which we will introduce in section 5.3 for assessing the plausibility of Assumption 5 and Assumption 6, with only a slight modification.

### v4

@@ 2.4   Interpreting the parallel trends assumption in the outcome
Thus far, we have formalized the target parameter and the identifying assumptions in $2\times 2$ DID-IV designs. Overall, in this design, we rely on the common DID identification strategy in both the denominator and the numerator of the Wald-DID estimand to identify the effects of the instrument on the treatment and the outcome. The interpretation of the parallel trends assumption in the outcome, however, is less clear than that of the parallel trends assumption in the canonical DID designs because it does not depend on the untreated outcome, but depends on the unexposed outcome.
@@
In this subsection, we clarify the interpretation of the parallel trends assumption in the outcome in $2\times 2$ DID-IV designs. To do so, we first uncover the selection mechanism behind the canonical DID-IV designs. Specifically, we investigate how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups. Then, we show that the average time trend in the outcome is equal to a weighted average of the indirect effect of time on the outcome through treatment. We call this an expected time gain and show that the parallel trends assumption in the outcome requires the homogeneous expected time gain between exposed and unexposed groups.
@@
First, we introduce the additional notation. Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$, which represents the treatment adoption process from period 00 to period 1 if the instrument path is $z=(0,0)$: the first element $D_{i,0}((0,0))$ represents the potential treatment choice without instrument in period 00 and the second element $D_{i,1}((0,0))$ represents the potential treatment choice without instrument in period $1$.
@@
Recall that in $2\times 2$ DID-IV designs, the units are allowed to take the treatment without instrument during the two periods. In this setting, one can see that time also plays the role of instrument; that is, time also affects one’s treatment choices during the two periods.
@@
Similar to the group variable $G_{i}^{Z}$ introduced in Subsection 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ as the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ as the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ as the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ as the time defiers. The time never-takers $NT^{T}$ and the time always-takers $AT^{T}$ do not change their treatment status without instrument during the two periods. In other words, their treatment choices are time-invariant in the absence of the policy shock (instrument). However, the time compliers $CM^{T}$ are induced to treatment in period $1$ without instrument, and the time defiers $DF^{T}$ leave the treatment in period $1$ without instrument; that is, their treatment choices are affected by time in the absence of a policy shock.
@@
By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs. Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, written in gray color in both tables.
@@
In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but not $D_{1}((0,0))$ for each unit (see Table 1). We use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$. In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but not $D_{1}((0,1))$ for each unit (see Table 2). If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time. If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument.
@@
We now consider the interpretation of the parallel trends assumption in the outcome. Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in the exposed and unexposed groups, respectively.
@@
Note that the parallel trends assumption in the outcome requires $\Delta_{0}=\Delta_{1}$. At the same time, $\Delta_{0}$ is estimable by the observed change in outcomes in an unexposed group during the two periods, whereas $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
To make the interpretations of $\Delta_{0}$ and $\Delta_{1}$ clear, we introduce the additional notation. For each time group $g^{t}\in\{AT^{T},NT^{T},CM^{T},DF^{T}\}$, let $w_{g^{t},e}\equiv Pr[G_{i}^{T}=g^{t}|E_{i}=e]$ denote the population share in group $E_{i}=e\in\{0,1\}$. Let $\Delta_{g^{t},e}$ denote the expectation of the potential outcome path in the absence of a policy shock for time group $g^{t}$ in group $E_{i}=e$:
@@
For instance, $w_{CM^{T},0}=Pr[CM^{T}|E_{i}=0]$ and $\Delta_{CM^{T},0}=E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=0]$ represent the population share and the expectation of the outcome path without instrument, respectively, for the time compliers in an unexposed group.
Each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are estimable in an unexposed group ($e=0$). We can identify the time group of unit $i$ in an unexposed group and can observe the evolution of the outcome without instrument. For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$ for unit $i$ in an unexposed group, we can figure out that unit $i$ is the time compliers $CM^{T}$, and can observe the potential outcome path $(Y_{i,0}(0),Y_{i,1}(1))$ without instrument for that unit. However, each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are unobservable in an exposed group ($e=1$) because we cannot specify the time group of unit $i$ in that group. For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$ for unit $i$ in an exposed group, we cannot distinguish whether unit $i$ belongs to the time compliers $CM^{T}$ or the time never takers $NT^{T}$ (see Table 1).
We can interpret each $\Delta_{g^{t},e}$ as the indirect effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$. First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes, respectively.
@@
This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods. If the exclusion restriction of time on the outcome holds, that is, $Y_{i,1}(1)=Y_{i,0}(1)$ and $Y_{i,1}(0)=Y_{i,0}(0)$ hold for all units $i$ (the stationary conditions hold), these quantities are equal to zero. Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms.
@@
where the weight $w_{g^{t},e}$ is non-negative and sums to one; that is, $\sum_{g^{t}}w_{g^{t},e}=1$ holds.
@@
Based on the above discussions, we conclude that the parallel trends assumption in the outcome (Assumption 6), that is, $\Delta_{0}=\Delta_{1}$ requires the same expected time gain between exposed and unexposed groups; the whole indirect effects of time on the outcome through treatment are the same on average between the two groups.
Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the indirect effect of the instrument on the outcome through treatment from that of time in an exposed group. The underlying idea to capture the former effect is parallel to that in DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period. In practice, if we have pre-exposed period data, we can indirectly check the validity of Assumption 6 by comparing the evolution of the outcome between exposed and unexposed groups. The formal discussion for the pre-trend test is provided in Subsection 5.3.
@@ 2.5   Relationship between DID and DID-IV
In this subsection, we describe the relationship between DID and DID-IV designs, and provide the view that DID-IV is DID with noncompliance of the assigned treatment.
Following Subsection 2.1, we introduce the notation in DID designs. Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program that distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$. In the canonical DID set up, $Z_{i}=(Z_{i,0},Z_{i,1})$ takes only two values: $Z_{i}=(0,1)$ if unit $i$ is assigned to the program in period $1$ (treatment group) and $Z_{i}=(0,0)$ if unit $i$ is not assigned to the program in period $1$ (control group). Let $E_{i}\in\{0,1\}$ denote the group variable: $E_{i}=1$ if unit $i$ belongs to the treatment group and $E_{i}=0$ if unit $i$ belongs to the control group. We refer to treatment and control groups as exposed and unexposed groups, respectively.
Note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units. First, we assume that all units are not allowed to receive the treatment in period 00.
@@
Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and that it should not receive the treatment in period $1$ if that unit belongs to an unexposed group.
@@
These restrictions imply that we have Tables 3-4 under DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
@@
Second, the target parameter in DID designs, the ATT is equal to the LATET in the canonical DID set up; this follows from Tables 3 - 4:
@@
Note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$). Therefore, we define the LATET as our target parameter in DID-IV designs in Subsection 2.2.
Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups. From Tables 3- 4, we can rewrite the parallel trends assumption in the untreated outcome as:
@@
where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively under DID designs.
@@
Overall, this subsection clarifies that DID-IV is a natural generalization of DID, and the differences between the two designs arise because DID-IV allows the noncompliance of the assigned treatment: $D_{i,t}\neq Z_{i,t}$. In the next section, we compare DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018).
@@ 2.6   Comparing DID-IV to Fuzzy DID
de Chaisemartin and D’Haultfœuille (2018) also investigate the identification assumptions for the Wald-DID estimand to capture some causal effects under the same setting as in this article. However, they formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID. The detailed discussion is given in Appendix E, and this section provides the overview of that discussion.
The main difference between dCDH and this article is in the definition of target parameter: we focus on the LATET, whereas dCDH focus on the SLATET (see Subsection 2.2). Accordingly, the identifying assumptions in Fuzzy DID designs are different from those in $2\times 2$ DID-IV designs. Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose the alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
In Appendix E, we first show that the identifying assumptions in Fuzzy DID designs ex-ante impose the strong restrictions on the treatment adoption behavior across units and these restrictions are asymmetric between exposed and unexposed groups.
Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Tables 5-6444Because dCDH consider repeated cross section settings, we can observe either $D_{0}(0)$ or $D_{1}(1)$ for each unit in Tables 5-6., which are in line with Tables 1-2. From these tables, one can see that dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group. Note that in our DID-IV design, we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by a monotonicity assumption. The difference between dCDH’s restrictions and ours is critical for empirical application; in contrast to the monotonicity assumption, which excludes the defiers in both groups, dCDH’s additional restrictions can be difficult to defend from the institutional knowledge of the problem in a given application. This is because these restrictions ex ante exclude the units who change the treatment adoption in the absence of a policy shock (instrument).
@@
Next, under Tables 5-6, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters. The one parameter measures the treatment effects among the type $CM^{Z}\land AT^{T}$, who are the sub-population of the compliers. The other causal parameter measures the treatment effects among the time compliers $CM^{T}$. The time compliers switch from non-treatment to treatment in period $1$ in the absence of the policy shock (instrument); that is, they are affected by time but not affected by the instrument. This decomposition result for the SLATET has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET. Second, the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher because this parameter is always contaminated by the treatment effects among the time compliers.
@@
Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided. During this discussion, we also clarify why dCDH ex ante exclude so many types in exposed and unexposed groups for the Wald-DID estimand to identify the SLATET.
@@
First, we show that dCDH’s argument regarding the use of the Wald-DID estimand is based on their reliance on the parallel trends assumption in DID designs. Recall that in our DID-IV design, we did not make the parallel trends assumption in the untreated outcome, but made the parallel trends assumption in the unexposed outcome. Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption is generally not sufficient to capture the average time trend of the outcome in an unexposed group. We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET. In the proof, we also demonstrate that imposing the parallel trends assumption in DID designs yields the strong restrictions on the treatment adoption behavior across units under Fuzzy DID designs.
@@
Next, after confirming that dCDH’s argument is based on the parallel trends assumption in DID designs, we argue that this assumption is unsuitable for DID-IV settings for three reasons. First, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group. Second, in most applications of the DID-IV method, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al. (2005)). For instance, in Black et al. (2005), this assumption requires the data to contain parents with zero education attainment, which is unrealistic in practice. Finally, in most of the DID-IV settings, we cannot test this assumption using pre-exposed period data, as some units can potentially adopt the treatment before period 00. We show that dCDH’s placebo test is incomplete for assessing the validity of this assumption. Surprisingly, we also show that their placebo test coincides with our pretrends test with a slight modification, introduced in section 5.3, to assess the plausibility of Assumption 5 and Assumption 6.